OpenStudy (anonymous):

PLEASE HELP ♥ The area of a rectangular swimming pool is 10x^2 – 19x – 15. The length of the pool is 5x + 3. What is the width of the pool?

OpenStudy (mertsj):

Area = length times width 10x^2-19x-15=(5x+3)(width)

OpenStudy (stamp):

Identify. A = Lw A = 10x^2 - 19x - 15 L = 5x + 3 Find w. A = Lw w = A/L Substitute. w = A/L w = (10x^2 - 19x - 15) / (5x + 3) Simplify.